Solve the equation logxe2e = eInx-e


1
Expert's answer
2021-10-26T15:00:08-0400

"2elne\/lnx=elnx-e"

"2=ln^2x-lnx"

"k=lnx"

"k^2-k-2=0"

"k_1=\frac{1+\sqrt{1+8}}{2}=2"

"k_2=\frac{1-\sqrt{1+8}}{2}=-1"


"lnx_1=2"

"x_1=e^2"


"lnx_2=-1"

"x_2=1\/e"